# WildAvatar: Web-scale In-the-wild Video Dataset for 3D Avatar Creation

### v2

@@ Appendix A Details in Data Pipeline
We provide more details on the data processing pipeline in addition to the main paper.
@@
Stage I: Human Bounding Box Detection and Tracking. We first obtain the bounding box of human subjects with an off-the-shelf state-of-the-art Yolo detection method [55]  [^2] The configuration of the detection model can be found here.. We only keep the video clip with at least one “person” instance with its detection threshold over $0.8$. The tracking step is based on finding the largest IOU overlay of bounding boxes among frames. We discard low-resolution human subjects whose bounding box areas are lower than $64\times 64$. To ensure the richness of the dataset, we only keep one “key subject” for each video, as clips of the same video may probably share the same key subject.
@@
Stage II: Human Segmentation Mask Extraction. To prepare key point exemplars for the Segment Anything, we first obtain the 2D keypoints $J_{2D}$ for human subjects using the popular HRNet [64]  [^3] The configuration of the 2D keypoint estimation model can be found here.. Given the 2D keypoint annotations, we are also able to discard over part-occluded subjects. Specially, we only keep the subject with the average confidence of 2D keypoints over $0.45$. For segmentation, we feed the 2D bounding box and the 2D keypoints into the sam_vit_h sub-model to extract the foreground mask.
Stage III: Coarse SMPL and Camera Estimation. We first estimate SMPL and camera parameters frame by frame, using a state-of-the-art single-image-based human pose and shape estimation method PARE [36]. To achieve better performance in complex scenes in the wild, we adapt the model pre-trained on the in-the-wild 3DPW dataset [68]. The PARE infers human body pose/shape parameters ($\theta$/$\beta$), and the global camera parameters (rotation matrix $R$ and the 3D offset $T$).
To retain the remaining video clips with considerable viewpoint shifts and human movements, we discard the clips with viewpoint angle changes lower than $\frac{\pi}{4}~{}rad$. And we automatically select $N=20$ most non-trivial frames, which keep the pose diversity to the greatest extent possible.
As mentioned in the main paper, we double-check the consistency between the SAM masks and the SMPL masks. Intuitively, the SMPL mask denotes the naked body, while the SAM mask contains the clothed body. Therefore, the SMPL mask should be mostly covered by the SAM mask (See Fig. A1 (a)  ~  (d)). We discard the subjects whose SAM masks are over 3×3$\times$3 × larger than their SMPL masks (See Fig. A1 (e)  ~  (h)). We also discard the subjects whose $10\%$ SMPL mask pixels are not covered by the SAM mask (See Fig. A1 (i)  ~  (l)).
@@
Stage IV: Refining SMPL and Camera In-the-loop. For high-quality annotations, here we refine the coarse SMPL parameters ($\theta,\beta$) and camera parameters ($R,T$) obtained in Stage I. To achieve temporally smooth results, we regularize the differences in parameters between adjacent frames, which is given by
$\mathcal{L}^{s}_{\theta}=\sum_{i=1}^{N-1}\left\|\theta^{i}-\theta^{i+1}\right\|_{2},\mathcal{L}^{s}_{R}=\sum_{i=1}^{N-1}\left\|R^{i}-R^{i+1}\right\|_{2},\mathcal{L}^{s}_{T}=\sum_{i=1}^{N-1}\left\|T^{i}-T^{i+1}\right\|_{2},$ (1)
$\mathcal{L}^{s}_{2D}=\sum_{i=1}^{N-1}\left\|\Pi(J_{3D}(\theta^{i},\beta^{i});R^{i},T^{i})-\Pi(J_{3D}(\theta^{i+1},\beta^{i+1});R^{i+1},T^{i+1})\right\|_{2},$
@@
Figure A1: SMPL and SAM consistency. The green/red borders denote good/bad outputs, respectively.
@@
$\mathcal{L}_{2D}=\sum_{i=1}^{N}\left\|\Pi(J_{3D}(\theta^{i},\beta^{i}))-J^{i}_{2D}\right\|_{2},\mathcal{L}_{\mathrm{mask}}=\sum_{i=1}^{N}d_{CD}(\Pi(V_{3D}(\theta^{i},\beta^{i})),M^{i}),$ (2)
where $d_{CD}$ denotes the chamfer distance between point clouds, and the $V_{3D}(\theta,\beta)$ infers the 3D vertex of the human body. The $M^{i}$ denotes the foreground d pixels cloud of the $i$ frame. We also regularize $\theta$ to avoid out-of-domain poses [7], using the Gaussian Mixture Model (GMM) prior [50]
@@
We empirically set the loss weights as $\lambda^{s}_{\theta}=100$, $\lambda^{s}_{R}=1000$, $\lambda^{s}_{T}=50$, $\lambda^{s}_{2D}=100$, $\lambda_{2D}=100$, $\lambda_{\mathrm{mask}}=100$ and $\lambda_{\mathrm{prior}}=0.1$. We adopt the LBFGS [61] optimizer with the learning rate $lr=1.0$. We optimize for $20$ steps on a single Geforce GTX 1080Ti GPU.
Qualitative comparisons between coarse and refined SMPL and camera parameters can be found in Fig. A2.
Figure A2: Comparison of the coarse and refined SMPL parameters.
Appendix B Additional Links
We release the WildAvatar dataset with a permanent doi 10.5281/zenodo.11526806. We also provide the Croissant metadata. The authors bear all responsibility in case of violation of rights, and confirm that this dataset is open-sourced under the CC BY 4.0 International license.

### v4

@@ Appendix B Details in Data Pipeline
In addition to the main paper, we provide more details on the data processing pipeline and filtering protocols.
@@
Stage I: Human Bounding Box Detection and Tracking. We first obtain the bounding box of human subjects with off-the-shelf state-of-the-art detection methods (e.g., Yolo [63] and Detectron2 [86]). We only keep the video clip with at least one “person” instance with its detection threshold over $0.8$ on all detection models. The tracking step is finding the largest IOU overlay of bounding boxes among frames. We discard low-resolution human subjects whose bounding box areas are lower than $64\times 64$. To ensure the richness of the dataset, we only keep one “key subject” for each video, as clips from the same video may probably share the same key subject.
@@
Stage II: Human Segmentation Mask Extraction. We first obtain the 2D keypoints $J_{2D}$ for human subjects using the popular HRNet [74] and DWPose [92]. Given the 2D keypoint annotations, we can also discard over part-occluded subjects. In particular, we only keep the subject with the average confidence of 2D keypoints over $0.65$. For segmentation, we feed the 2D bounding box and the 2D keypoints into the sam_vit_h sub-model to extract the foreground mask.
Stage III: Coarse SMPL and Camera Estimation. We first estimate SMPL and camera parameters frame by frame, using state-of-the-art single-image-based human pose and shape (HPS) estimation methods [43, 23]. To perform better in complex scenes in the wild, we adapt the model pre-trained on the in-the-wild 3DPW dataset [79]. The HPS models infer human body pose/shape parameters ($\theta$/$\beta$) and the global camera parameters (rotation matrix $R$ and the 3D offset $T$). To retain the remaining video clips with considerable viewpoint shifts and human movements, we discard the clips with viewpoint angle changes lower than $\frac{\pi}{4}~{}rad$. We also automatically select the most non-trivial $N=20$ frames, which keeps the pose and viewpoint diversity to the greatest extent possible. As mentioned in the main paper, we double-check the consistency of the SAM and SMPL masks. Intuitively, the SMPL mask denotes the naked body, while the SAM mask contains the clothed body. Therefore, the SMPL mask should be mostly covered by the SAM mask (See Fig. F (a)  ~  (d)). We discard the subjects whose SAM masks are over 3×3$\times$3 × larger than their SMPL masks (See Fig. F (e)  ~  (h)). We also discard the subjects whose $10\%$ SMPL mask pixels from main bodies are not covered by the SAM mask (See Fig. F (i)  ~  (l)). Similarly, we double-check the consistency of the 2D keypoints from 2D pose and SMPL estimations and discard the clips with the averaged PCK less than $0.85$.
@@
Stage IV: Refining SMPL and Camera In-the-loop. We refine the coarse SMPL parameters ($\theta,\beta$) and camera parameters ($R,T$) obtained in Stage I for high-quality annotations. To achieve temporally smooth results, we regularize the differences in parameters between adjacent frames, which is given by
$\mathcal{L}^{s}_{\theta}$ $=\sum_{i=1}^{N-1}\left\|\theta^{i}-\theta^{i+1}\right\|_{2},$ (1)
$\mathcal{L}^{s}_{R}$ $=\sum_{i=1}^{N-1}\left\|R^{i}-R^{i+1}\right\|_{2},$
$\mathcal{L}^{s}_{T}$ $=\sum_{i=1}^{N-1}\left\|T^{i}-T^{i+1}\right\|_{2},$
$\mathcal{L}^{s}_{2D}$ $=\sum_{i=1}^{N-1}\left\|\Pi(J_{3D}(\theta^{i},\beta^{i});R^{i},T^{i})\right.$
$\quad\left.-\Pi(J_{3D}(\theta^{i+1},\beta^{i+1});R^{i+1},T^{i+1})\right\|_{2},$
@@
$\mathcal{L}_{2D}$ $=\sum_{i=1}^{N}\left\|\Pi(J_{3D}(\theta^{i},\beta^{i}))-J^{i}_{2D}\right\|_{2},$ (2)
$\mathcal{L}_{\mathrm{mask}}$ $=\sum_{i=1}^{N}(M_{rd}(\theta^{i},\beta^{i})),M^{i}),$
where $M_{rd}(\theta,\beta)$ denotes the rendered mask related to the SMPL parameters. The $M^{i}$ denotes the foreground of the $i$ frame. We also regularize $\theta$ to avoid out-of-domain poses [7], using the Gaussian Mixture Model (GMM) prior [58]
@@
We empirically set the loss weights as $\lambda^{s}_{\theta}=100$, $\lambda^{s}_{R}=1000$, $\lambda^{s}_{T}=50$, $\lambda^{s}_{2D}=100$, $\lambda_{2D}=100$, $\lambda_{\mathrm{mask}}=100$ and $\lambda_{\mathrm{prior}}=0.1$. We adopt the LBFGS [69] optimizer with the learning rate $lr=1.0$. Qualitative comparisons between coarse and refined SMPL and camera parameters can be found in Fig. E. Stage IV boosts local alignments on extremities.
Figure B: Visualization of SMPL overlay on unusual camera viewpoints. Our pipeline can show robust annotations on unusual camera viewpoints and body poses.
Efficiency. Data collection and annotation efficiency are also crucial for data scale-up and application. Despite the complex multiple-stage design, our data-collecting pipeline only takes less than $200$ seconds to generate full annotations for a $20$ seconds in-the-wild video clip.
Appendix C Details in Pipeline Evaluation
Dataset. We evaluate our four-stage pipeline on the in-the-wild EMDB [40], which is widely recognized for its challenging and diverse real-world scenarios. We use the EMDB-1 split to evaluate the camera-coordinate performance. EMDB-1 contains $17$ sequences totaling $13.5$ minutes.
Metrics. For joints, we compute error on the $24$ main joints of the human body under the SMPL convention. As for vertex, we calculate the point-to-point corresponding error on the SMPL vertices. We report quantitative results on MPJPE, PA-MPJPE [33], and PVE [89]. MPJPE (Mean Per Joint Position Error) calculates the mean distances between the predicted and ground-truth 3D joints after the translation alignment at the pelvis joint. The predicted or ground-truth 3D joints are regressed from corresponding pose and shape parameters. PA-MPJPE (Procrustes analysis MPJPE) calculates the mean distances between the predicted and ground-truth 3D joints after Procrustes Analysis [24], including alignment in scale, translation and rotation. PA-MPJPE mainly focuses on the quality of pose and shape estimation, regardless of global rotation. PVE (Per Vertex Error) calculates the mean distances between the vertices on the human mesh without any alignment, which evaluates the reconstruction accuracy of the human surface.
Qualitative Comparisons. Qualitative comparisons on the EMDB dataset are also shown in Fig. A (a) ~ (d). Compared to previous state-of-the-art HMR2.0 [23], our pipeline can adapt to complex environmental scenes and output reasonable results in uncommon scenarios. For example, our pipeline can accurately predict 1) the foot and ankle pose of Fig. A (a) & (c); 2) the head and neck pose of Fig. A (b) & (d); 3) the global alignment of Fig. A (a), (b) & (d). Qualitative comparisons on the 3DPW dataset are also shown in Fig. A (e) ~ (h). Compared to the official ground truth (GT) from expensive IMUs, our annotations also exhibit better alignment on 1) the foot and ankle pose of Fig. A (e) ~ (g); 2) the head and neck pose of Fig. A (g); 3) the hand pose of Fig. A (h). These comparisons validate the annotation quality of our pipeline for in-the-wild videos.
Appendix D License, Statistics and Visualizations
The authors bear all responsibility in case of violation of rights and confirm that this dataset is open-sourced under the S-Lab License 1.0 license. We shall enforce strict regulations when applying our code and data to mitigate potential negative social impacts. $69.1\%$ / $30.9\%$ scenes in WildAvatar are indoor/outdoor, respectively. $45.3\%$ / $54.7\%$ of the scenes in WildAvatar have single/multiple human(s), respectively. And $34.6\%$ / $65.4\%$ subjects in WildAvatar are male/female, respectively. More visualization of SMPL overlay on unusual camera viewpoints can be found in Fig. B. More RGB examples from the proposed WildAvatar dataset can be found in Fig. C, and examples of different SSIOU ranges can be found in Fig. D.
Figure C: More RGB examples from the proposed WildAvatar dataset. The best view zoomed in on-screen for details.
Figure D: Examples of different SSIOU ranges. SSIOU rises from up to down. The last row shows the SSIOU larger than $1.9$. The samples with SSIOU larger than $1.9$ are mostly caused by loose dresses rather than erroneous SMPL fitting.
Figure E: Comparison of the coarse and refined SMPL parameters. The coarse SMPL annotations are from Stage III, and are later refined in Stage IV. The refined SMPL parameters achieve better alignments to the raw RGB images.
Figure F: SMPL and SAM consistency. The green/red borders denote good/bad outputs, respectively. Note that the SMPL annotations are coarse results from Stage III, which are later refined in Stage IV.